A 50 mm diameter stream of water strikes a 0.5 m by 0.5 m square plate which is at an angle of 30 degrees with the stream’s direction. The velocity of the water in the stream is 20 m/s and the jet strikes the door at its centre of gravity. If the plate is hinged at A, by neglecting the friction and self-weight of the plate, determine the normal force P applied at the upper edge of the plate to keep the plate in equilibrium. You may assume that the velocities of the jet remain unchanged after split.

Please help me
1
Expert's answer
2019-04-11T09:54:33-0400


Newton’s second law:

"F_w \Delta t = m(v-0)\\"

"F_w\Delta t = \rho V v"

"F_w\Delta t = \rho S v \Delta t v"

"F_w = \rho\frac{\pi d^2}{4} v^2 = 1000 \frac{kg}{m^3} \frac{\pi \cdot 25 \cdot 10^{-4}m^2}{4} \cdot 400 \frac{m^2}{s^2} = 250\pi N"


"P = \sin (\phi) F_w = sin(30^\circ) F_w = 250\pi \cdot 0.5N = 125\pi N = 392.7N"


Answer: 392.7N